### v3

# In-Context Reinforcement Learning for Variable Action Spaces

Abstract
Recently, it has been shown that transformers pre-trained on diverse datasets with multi-episode contexts can generalize to new reinforcement learning tasks in-context. A key limitation of previously proposed models is their reliance on a predefined action space size and structure. The introduction of a new action space often requires data re-collection and model re-training, which can be costly for some applications. In our work, we show that it is possible to mitigate this issue by proposing the Headless-AD model that, despite being trained only once, is capable of generalizing to discrete action spaces of variable size, semantic content and order. By experimenting with Bernoulli and contextual bandits, as well as a gridworld environment, we show that Headless-AD exhibits significant capability to generalize to action spaces it has never encountered, even outperforming specialized models trained for a specific set of actions on several environment configurations.
Machine Learning, ICML
1 Introduction
Figure 1: Variable Action Spaces: We consider four types of novel action spaces different from the one used during training. Permuted Train Actions maintains the action set contents but reorders its elements. Test Actions introduces a completely new action set with an increased size. It is important to consider that some models may be architecturally limited to a fixed action set size. To evaluate the performance of such models on unseen actions, we adjust the size of a new set to be compatible with the model output. Therefore, we slice the first actions from the Test Actions set. Lastly, a new action space might include both the seen Train and unseen Test actions, depicted as the All Actions set.
Figure 2: Headless-AD Architecture: Compared to AD, Headless-AD introduces four new components. (1) We remove the output linear head, making the model directly predict the action embedding. That allows us to avoid a direct connection between the model and action space size, contents and ordering. (2.1) At each training step, we generate random action embeddings for each action in the action set. (2.2) We convert actions in the context into their embeddings and pass them as the model input. This prepares the model for unseen actions, forcing it to infer action semantics from the context. (3) As the model loses prior knowledge about action space structure, we pass the generated action embeddings as a prompt to aid the model in sensible action selection. (4) We convert a prediction vector into a distribution over actions based on the similarities between the prediction and previously generated action embeddings. To increase the probability of correct actions, we use contrastive loss instead of cross-entropy.
Figure 3: Algorithm Distillation Struggles with Novel Action Spaces: Despite its good results on the train action set, AD’s performance diminishes when the action semantics change, either due to a permutation or substitution. It is important to note that augmenting the training data with permuted action sets does not lead to increased performance, signifying that action set invariance should be enforced from a model design standpoint. Additionally, it is impossible to apply a trained AD model to a larger action set. On the graph, the bars are the success rate values on the gridworld environment (described in Section 4.3) obtained after evaluating each of the action sets visualized in Figure 1, averaged over 5 runs. Altered Semantics aggregate the values from the Permuted Train Actions and Sliced Test Actions sets. Altered Size aggregates the values from Test Actions and All Actions. See Section 4.3 for more information about the construction of the action sets.
The transformer architecture, first introduced by Vaswani et al. (2017), has been widely adopted in key areas of machine learning, including natural language processing (Radford et al., 2018; Devlin et al., 2018), computer vision (Dosovitskiy et al., 2020) and sequential decision-making (Chen et al., 2021). One major feature of transformers is in-context learning (ICL), which makes it possible for them to adapt to new tasks after extensive pre-training (Brown et al., 2020; Liu et al., 2023). Recent developments, such as Algorithm Distillation (AD) by Laskin et al. (2022) and Decision Pretrained Transformer (DPT) by Lee et al. (2023), have successfully employed transformer ICL abilities in sequential decision-making. These models are capable of predicting the next action based on a query state and history of environment interactions, which inform them about task objectives and environment dynamics. While effective at generalizing across various reward distributions (Laskin et al., 2022; Lee et al., 2023) and transition functions (Raparthy et al., 2023), their adaptability to new action spaces remains unexplored and limited by architectural constraints.
Creating models that can adapt to new action spaces is essential for building the foundation of decision-making systems in order to enable large-scale pretraining across various environments and address real-world problems (Jain et al., 2020; Chandak et al., 2020; London & Joachims, 2020; Jain et al., 2021). With this in mind, our research focuses on variable discrete action spaces, with the notion of variability illustrated in the Figure 1. In our study, we reveal the limitations of the Algorithm Distillation (AD) model from prior work, such as its diminished performance upon changes in action semantics as well as architectural constraints when handling varying action space sizes (see Figure 3).
Our solution, Headless-AD, is an architecture and training methodology tailored to effective generalization on new action spaces. We employ an approach similar to Wolterpinger (Dulac-Arnold et al., 2015) and Headless-LLM (Godey et al., 2023) by encoding actions with random embeddings (Kirsch et al., 2023) and directly predicting these embeddings. This way, we remove the direct connection between the model output layer and the action space structure.
Through experiments using Bernoulli and contextual bandits, and a darkroom environment with changing action spaces, we demonstrate that Headless-AD is capable of matching the performance of the original data generation algorithm and scaling to action spaces up to 5x larger than those seen during training. We also observed that Headless-AD can even outperform AD when they are both trained for the same action space, especially when evaluated on larger action sets. To summarize, our contributions are as follows:
  • We show that AD struggles with generalization on novel action spaces (Section 2).
  • We extend AD with a modified model architecture and a training strategy, called Headless-AD, for it to acquire the ability to adapt to new discrete action spaces (Section 3). We demonstrate the strong generalization capabilities of Headless-AD on Bernoulli and contextual bandits, and darkroom environments (Section 4).
  • We perform ablations on the loss and the prompt format to highlight the importance of Headless-AD’s design choices (Section 5).
Figure 4: Algorithm Regret under Variable Reward Distributions in Bernoulli Bandit. The graph compares regret for Random, Thompson Sampling, and Headless-AD across distinct reward distributions in the Bernoulli Bandit environment, averaged from five seeds. During training, the high reward was $95\%$ more likely to distribute across the odd arms. During testing, it either switched to the even arms or a uniform distribution. Note that Headless-AD maintains high performance in all configurations, proving its ICL capabilities at generalizing to novel tasks, represented by changes in reward distribution. Data is aggregated from bandit problems with $4-20$ arms, reflecting the training conditions.
2 Algorithm Distillation Struggles with Novel Action Spaces
Algorithm Distillation (AD) (Laskin et al., 2022) is a transformer model trained to autoregressively predict the next action given the history of previous environment interactions and a current state. Formally, the history is defined as:
$h_{t}=\left(o_{0},a_{0},r_{0},\dots,o_{t},a_{t},r_{t}\right),$
where $o$ are the observations, $a$ are the actions and $r$ are the rewards. The probabilities of each action are given by a model $P_{\theta}(A=a^{(n)}_{t}|h_{t-1},o_{t})$, where $n$ is the action index, $t$ is a timestamp and $\theta$ are the model weights. AD is pretrained on data logged by a training agent, and the context size should be sufficiently large to span multiple episodes in order to capture policy improvement. This way, AD learns an improvement operator that increases performance entirely in-context when applied to novel tasks.
The model output is a probability distribution across the action set, derived from a linear projection and a softmax function. As highlighted in Figure 3, this structure causes a fundamental limitation in AD’s adaptability to new action spaces, as the output dimension is predetermined. To accommodate an action set of a different size than the one used during training, the model’s final layer must be redefined and the model retrained. Moreover, even with a constant action space size, the model’s efficacy diminishes if the action semantics are altered. The reason for this is the classifier nature of the model, which associates each dimension with a particular meaning of an action. Since augmenting the dataset with permuted action sets does not lead to improvement, it signifies that action set invariance should be enforced from a model design standpoint.
Figure 5: Algorithm Regret under Increasing Amount of Arms in Bernoulli Bandit. This series of plots shows the regret of Thompson Sampling, AD, and Headless-AD algorithms over evaluation steps in environments with $20-50$ arms, averaged from five seeds with $100$ bandits each. Although Headless-AD has been trained on bandits with up to $20$ arms, it performs well, matching or outperforming other algorithms in larger arm settings without additional training. Note that AD was retrained from scratch for each task with a different number of arms.
3 Headless-AD
To mitigate the action space limitations of AD, we propose Headless-AD, a new architecture that improves on AD by omitting the final linear layer and incorporating three key modifications. The Headless-AD architecture and data flow is visualized in Figure 2.
Direct Prediction of Action Embeddings: The model output is modified to yield an action embedding rather than a probability distribution over actions. This alteration makes the model independent of action set size and order, granting it permutation invariance. The InfoNCE Contrastive Loss (Oord et al., 2018), diverging from its usual role in representation learning (Jaiswal et al., 2020), serves as a regression objective to reinforce the similarity between the model prediction and the subsequent action in the data. All other actions are treated as negative samples:
$L=-\mathop{\mathbb{E}}\left[{\log{\frac{f(\hat{a}_{t},a_{t})/\tau}{\sum_{a\in A}{f(\hat{a}_{t},a})/\tau}}}\right],$
where $\tau$ is a temperature parameter. We used dot-product as the similarity function $f$.
Random Action Embedding Strategy: To prepare the model for potential new actions during testing, we transition from static learned action embeddings to dynamic embeddings generated by a mapping function $g:~{}\mathbb{A}~{}\rightarrow~{}\mathbb{R}^{n}$, which produces a unique random encoding for each action in a batch at the start of every training step. The mapping is shared across all batch instances and is consistent along the context sequence, i.e., actions with index $i$ in their respective action sets will all be mapped to the same embedding. The random embeddings are constrained as orthogonal vectors within a unit sphere. The details of the sampling process can be found in Appendix F. This approach makes the model deduce action semantics from the interaction history. The orthogonality also clarifies the interpretation of similarity measures as logits because it isolates the similarity of each action, thus preventing overlap in the probability distribution. Moreover, employing random embeddings enhances data variety, which has been demonstrated to boost in-context learning for RL agents (Kirsch et al., 2023; Lu et al., 2023).
Action Set Prompt: To address the model’s lack of awareness of the action space structure, we prepend the input with a sequence of embeddings for all available actions.
The modified input format is thus represented as:
$h_{t}=\left(a^{emb,0},\dots,a^{emb,N},o_{<t},a^{emb}_{<t},r_{<t},o_{t}\right),$
where $N$ is the action set size and $a^{emb}=g(a)$.
We employ two methods to select actions during inference: (1) nearest neighbor selection, and (2) probabilistic sampling based on the similarity to the predicted action embedding. We treat the specific choice of method as a hyperparameter. An illustrative code snippet with Headless-AD’s training procedure can be found in Appendix I.
4 Experiments
As Headless-AD extends and improves on AD, we check it in two different aspects. Firstly, it should maintain In-Context Learning abilities and thus generalize well to new tasks. Secondly, it should show high performance on action spaces different from the one seen during training. All of the following environments are designed specifically to check both of the above aspects. In our experiments, we use the TinyLLaMa (Zhang et al., 2024) implementation of the transformer model and AdamW optimizer (Loshchilov & Hutter, 2017). All environment specific hyperparameters are listed in Appendix H.
4.1 Bandit
Figure 6: Contextual Bandit Regret Comparison: The Train set consists of trajectories of LinUCB trained for $300$ steps on contextual bandits with $4-20$ arms. All models are also evaluated for $300$ steps. The results are averaged over $5$ seeds, each seed containing $100$ environments. AD requires retraining on each new action set and, while showing good performance on lower space sizes, it fails to converge on larger ones. Due to its variable-size action sets, AD was not tested on the Train set. Conversely, Headless-AD, trained exclusively on the Train set, is successful at both learning effectively within this environment and generalizing to new action sets.
Motivation: This experiment checks Headless-AD’s abilities on a toy task, where the environment does not have a notion of state and returns binary rewards.
Setup 1: In our first experiment, we examined the model’s robustness to distributional shifts in rewards. We used a Bernoulli bandit where each arm is associated with a mean $\mu$ and the reward after pulling an arm $i$ is generated by $\text{Bernoulli}(\mu)$. The training dataset consists of bandits with $4-20$ arms, i.e., different action set sizes. Additionally, to evaluate the ICL capabilities of models, the reward distribution over the arms differed between train and test (Laskin et al., 2022). Specifically, the training data consisted of bandits where in $95\%$ of cases the odd-numbered arms were assigned a random $\mu\in[0.5,1]$ and even-numbered arms received $\mu\in[0,0.5]$. For other $5\%$ of bandits, the ranges were swapped between odd and even arms. The test distribution also consisted of bandits with $4-20$ arms, but the reward distribution was switched either to even arms or was uniform, i.e., all arms were assigned $\mu\in[0,1]$. Learning histories were generated using Thompson Sampling algorithm for a total of $10,000$ bandit instances with $300$ steps each. The evaluation was performed on $100$ bandits in each reward distribution, included $5$ seeds, and the algorithms were rolled out for $300$ steps.
Results and Discussion 1: As depicted in Figure 4, the Headless-AD model demonstrated strong generalization capabilities by almost reaching the performance results set by the traditional Thompson Sampling algorithm under each test distribution.
Setup 2: In our second experiment, we evaluated the transferability of Headless-AD to new action set sizes. Training distribution remained the same as in the previous experiment. Each evaluation dataset consisted of a fixed amount of $20$, $25$, $30$, $40$ and $50$ arms and a uniform distribution of rewards. AD was trained on fixed-size bandits corresponding to each evaluation dataset. The training and evaluation reward distributions are the same as for Headless-AD.
Results and Discussion 2: Figure 5 showcases that Headless-AD can effectively maintain its performance as the action space grows, without necessitating any retraining. Moreover, Headless-AD even outperforms specially trained AD, especially on larger action sets. Note that Headless-AD’s performance curves resemble TS’s performance curves, signifying that it has indeed learned a policy improvement operator generic enough to apply to unseen action set sizes.
4.2 Contextual Bandit
Motivation: To validate the sustained performance of Headless-AD, we progress to a more complex Multi-Armed Bandit (MAB) extension that integrates states and real-valued rewards.
Setup: Each time step presents a context with arm features modeled as two-dimensional vectors, and rewards are generated with a standard deviation $\sigma=1$. The data is created using a LinUCB (Li et al., 2010) algorithm trained over $300$ steps, on bandits with $4-20$ arms. We used $100$ contextual bandits, $5$ seeds and $300$ evaluation steps.
Results and Discussion: As shown in Figure 6, Headless-AD’s performance is on par with LinUCB across varied arm counts, while also being capable of adept generalization to new environments. While AD also reaches the performance of LinUCB on lower space sizes, it has problems converging on larger action spaces sizes, in addition to requiring a specific retraining. This serves to highlight the advantages of Headless-AD for use in even more complex environments than toy Bernoulli bandits.
Figure 7: Darkroom Environment Success Rate: The chart displays mean success rates and their standard deviations from five training seeds, comparing performance in fixed and variable action spaces, along with the models’ adaptability to new goals. Train Actions refers to the fixed-size action set used for training. Test Actions include exclusively unseen actions, while All Actions combine both Train and Test, with set sizes expanded to $75$ and $125$ respectively. Since the output dimension changes, AD requires retraining. To assess AD’s adaptability to the changed action semantics, we either permute Train Actions or replace it with the first $50$ actions from the Test set. In this case, AD does not require retraining, but it still exhibits diminished performance. Conversely, Headless-AD, while trained solely on Train Actions, delivers strong and stable performance across all action set variants, and even surpasses specially trained AD on larger set sizes.
4.3 Darkroom
Motivation: In this experiment, we delve into a more sophisticated Markov Decision Process (MDP) framework, constructing five distinct action spaces aimed at demonstrating the architectural constraints of AD. We then show how Headless-AD’s architecture is engineered to navigate the complexities presented by each of these diverse action sets.
Setup: The Darkroom environment, inspired by Chevalier-Boisvert et al. (2018) and Jain et al. (2020), consists of a $N\times N$ grid where the agent needs to reach a specific cell for a reward. Action space consists of $3$-step sequences of $5$ atomic actions: up, down, left, right, noop. As a result, the environment offers $5^{3}$ possible actions. The agent earns a reward of $1$ if the trajectory induced by the action sequence passes through a goal cell, after which the episode finishes. Otherwise, the reward is 00. As an observation, the environment offers only the current coordinates of the agent, so the goal information can only be obtained from the agent’s memory. We divide the goals into disjointed sets used for training and testing in order to evaluate Headless-AD’s in-context learning abilities. Furthermore, the action set is randomly split into train and test sets, each including $50$ and $75$ actions respectively, to create five distinct spaces for assessing various generalization aspects. These action sets are visualized in Figure 1.
Train Actions. Comprising the training split, this set represents the actions encountered during model training.
Test Actions. Comprising the test split, this set assesses Headless-AD’s generalization on novel and larger action sets.
All Actions. Combining both training and testing actions, this set contains $125$ actions and is $2.5$ times larger than the training set. Its aim is to challenge Headless-AD to effectively integrate seen and unseen actions.
Permuted Train Actions. Shuffled training set, meant to test the model’s adaptability to reordered action spaces that are comprised of the same actions.
Sliced Test Actions. Tailored to match the training set size, this set contains a slice of the first $50$ actions from the test set. Its aim is to check the models’ generalization abilities on unseen actions while maintaining the action set size.
In scenarios where the action space exceeds the size of the training set, we analyze the performance of an AD model retrained from scratch. The data generation algorithm is Q-learning, executed over $200$ episodes for each environment. Further details on model hyperparameters are available in Appendix H.
Figure 8: Impact of Action Set Prompt Ablation: (a) The inclusion or exclusion of the prompt, which lists all possible action embeddings, does not substantially affect the number of distinct actions explored by the agents. This plot shows an averaged number of actions values across all action sets. (b) The exclusion of the prompt significantly impacts performance in environments like the Bernoulli Bandit, while the effect is less prominent in the Darkroom environment. This plot shows average performance values across all action sets. Performance on the y-axis represents the normalized regret on the Bernoulli Bandit environment, and the success rate on Darkroom. (c) The exclusion of the prompt has an impact on performance that becomes increasingly significant with the growth of the action space. The trend suggests that, as the number of potential actions rises, the ability to strategically sample actions based on the action set prompt becomes increasingly vital for model efficacy.
Results and Discussion: Figure 7 illustrates the generalization abilities of both AD and Headless-AD. First, note that both models maintain their performance when transitioning to novel tasks, thus fulfilling their purpose as ICL-RL models. However, this environment is mainly designed to challenge the models’ generalization abilities on novel action spaces. Following the scope of the action space novelty we set earlier, we study the models’ ability to address changed action semantics and variable action set sizes.
While AD achieves high performance on the train action set, its limitations become evident when the action space is changed. The first limitation is AD’s action set size constraint. The linear layer at the end of its network fixes the size of the output dimension, something that cannot be modified once the model is trained. Increasing the amount of options, as was done in the Test Actions and All Actions sets, requires reinitializing the output dimension and thus retraining the model, which demands additional time and resources. In contrast, once trained, Headless-AD swiftly adapts to changes in the action set size without losing performance.
The second limitation is AD’s reliance on a stationary action space structure. Due to the classifier nature of AD’s network, it learns to associate each output dimension with a specific action meaning. When action semantics change either due to a permutation of a seen action, as in Permuted Train Actions, or due to a substitution of completely new actions, as in Sliced Test Actions, AD’s performance degrades. We checked whether this problem may be solved by biasing the training distribution to have a structure resembling the one during testing. We trained AD on permuted action sets while leaving the data and the model architecture the same. However, as one can see in Figure 3, this training procedure results only in a slightly decreased performance of AD in the Permuted Train setting. An additional graph depicting the performance on each of the (action set, goal type) pairs can be found in Appendix C. Meanwhile, Permuted Train Actions does not pose a challenge for Headless-AD as, by design, it is invariant to specific action order. Most importantly, Headless-AD maintains its performance even when completely new actions are introduced, despite having never seen them in training. We attribute this feature to our success in making the model infer action meaning from the context.
5 Ablations
Action Set Prompt: We assess the impact of omitting the action embedding enumeration from the model context. The tuned hyperparameters can be found in Table 4.
Figure 7(a) reveals that removing the action set prompt does not significantly alter the number of unique actions attempted by each model. This suggests that models without the prompt continue to sample a diverse range of actions, comparable to the original setup. However, as shown in Figure 7(b), the presence of the prompt does affect performance. We hypothesize that, without the action set prompt, the model lacks explicit knowledge of the action space, which may lead to a more randomized sampling within the embedding space. Conversely, with the prompt in place, the model can directly target specific action embeddings.
Figure 7(b) shows that the absence of the prompt is more detrimental in environments such as the Bernoulli Bandit, where each decision has a direct impact on the final outcome due to the environment’s single-episode structure. However, in multi-episode environments such as the Darkroom, the early lack of action space information becomes less impactful over time, as the model eventually encounters all actions through the context. This divergence highlights the prompt’s critical role in enabling the model to make informed choices in environments where each option has immediate and lasting consequences. Note that, as one can see in Figure 7(c), the impact of action set prompt ablation on the performance gets more pronounced as the number of arms grows.
Contrastive Loss: In principle, the prediction of action embedding can be facilitated by various ways. Here, we consider the performance of a model that is asked to directly copy the corresponding embedding from its context to the output. To achieve it, we use mean squared error (MSE) loss instead of contrastive loss. In this case, probabilistic interpretation becomes less meaningful, which is why the nearest neighbor of the predicted vector is chosen as an action index, without sampling. The tuned hyperparameters can be found in Table 5.
Figure 9: Evaluating the Effectiveness of Contrastive vs. MSE Loss: This figure presents a comparative analysis of contrastive loss and mean squared error (MSE) loss on the training efficiency and performance metrics of our model. All results are averaged across various action sets within their respective environments and are based on data from five distinct training seeds. Performance on the y-axis represents the normalized regret for the Bernoulli Bandit environment, and the success rate for Darkroom. (a) depicts the impact on the model’s ability to explore: in the Bernoulli Bandit environment, the model trained with MSE loss exhibits a limited exploration range, focusing on only a small subset of the available actions. Conversely, on the Darkroom environment there is a slight increase in the number of actions attempted. Regardless of how the MSE loss affects the amount of actions used, the final performance of the model on both environments drops significantly.
Figure 8(a) illustrates a significant decline in the variety of actions attempted by the model under this configuration within the bandit environment, showing that the model concentrates only on a fraction of the action set. Though the final regrets for this model are far from random, as depicted in Figure 8(b), we point out that this result is because the model skips exploration phase and goes directly to an exploitation of suboptimal actions. That allows it not to waste time on even more suboptimal ones via exploration. Appendix D shows the in-context curves of both models, providing more insight into MSE’s effect on model quality.
Conversely, on the Darkroom environment, the loss substitution leads to a marginal increase in the amount of attempted actions. However, as one can see in Figure 8(b), this does not lead to increased performance. In fact, the performance of this model on the Darkroom environment is similar to random behavior.
In summary, although the employed neural network is capable of learning an improvement operator and demonstrating ICL capabilities, the implemented loss function plays a crucial role in the success of this learning process. We emphasize the importance of our design choice to use contrastive loss by showing that a more naive approach of directly copying the action embeddings, as incentivized by MSE loss, results in underperformance.
6 Related Work
Here, we discuss previous research in adapting RL to environments with variable action spaces. For an extended Literature Review, see Appendix B.
Recent research by Chandak et al. (2020); Ye et al. (2023) has focused on scenarios where the amount of available actions grows during evaluation by introducing new actions to an existing set. However, their model requires fine-tuning when new actions are introduced, while our model does not require parameter updates. Additionally, our research explores a broader spectrum of dynamic action spaces, encompassing the addition, removal, and substitution of actions. Lastly, we work in a setting where the action set remains constant throughout model evaluation.
Kirsch et al. (2022) present the concept of SymLA, a methodology that ensures resilience to changes in input and output sizes and permutations. This is achieved by integrating symmetries into a neural network model by representing each neuron with uniformly structured RNNs and data flow with message-passing connections. However, we suggest a simpler approach by extending the well-known transformer architecture. Additionally, we demonstrate the high performance of our model on more complex environments.
Further developments by Jain et al. (2020) include a specific module designed to generate action representations informed by observed trajectories after action application. Additionally, Jain et al. (2021) delve into variable action spaces, placing a specific emphasis on modeling the interconnections among actions to improve the quality of action representations. Our work, however, adopts a more implicit approach to inferring action-related information.
In a similar work, Lu et al. (2023); Kirsch et al. (2023) employ random projections for action encoding, a technique we also utilize. This method facilitates training across multiple domains with actions of varying sizes. Nonetheless, their focus is predominantly on continuous spaces, while our focus is on discrete action spaces.
To the best of our knowledge, Headless-AD is the first study to explore variable discrete action spaces for in-context reinforcement learning.
7 Conclusion
In our work, we have introduced a new architecture that extends Algorithm Distillation (AD) for environments with variable action spaces, achieving invariance to their structure and size. Our approach consists of discarding the last linear layer, granting invariance to the action space structure, and making the model infer action semantics from the context, preparing it for the introduction of novel actions. We have demonstrated Headless-AD’s capability to generalize across new action spaces on a set of environments. We also observed its performance gains over vanilla AD, especially on larger action spaces. We hypothesize that this is due to the augmentation of the dataset by random embeddings, which was shown to improve the generalization abilities of agents (Kirsch et al., 2023). Headless-AD marks the progress toward versatile foundational models in RL, ones capable of operating across an expanded range of environments. We hope that Headless-AD inspires further development of models that can adapt to any action space beyond discrete ones.
8 Broader Impact
The work presented in this paper is aimed at advancing the field of Machine Learning. There can be many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Background
Partially Observable Markov Decision Process. A Markov Decision Process (MDP) is defined by a tuple $(S,A,P,R)$, where $s_{t}\in S$ denotes a state, $a_{t}\in A$ an action, $p(s^{\prime}|s_{t}=s,a_{t}=a)$ the transition probability from state $s_{t}$ to $s_{t+1}$ after taking action $a_{t}$, and $R(s,a)$ the reward for action $a$ in state $s$ (Sutton & Barto, 2018). An agent $\pi$ observes the state $s_{t}$, selects action $a_{t}\sim\pi(\cdot|s_{t})$, and receives the subsequent state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$ and reward $R(s_{t},a_{t})$. In POMDPs, the agent receives an observation $o_{t}$ instead of the full state $s$, which contains partial information about the MDP’s real state. In the context of our work, $o_{t}$ may lack goal information, requiring inference from the agent’s memory.
Multi-Armed Bandits. A Bernoulli multi-armed bandit (MAB) environment consists of $N$ arms $a_{i}\in A$, each associated with a mean $\mu_{i}$ (Sutton & Barto, 2018). Pulling an arm $a_{i}$ yields a reward $r_{i}\sim\text{Bernoulli}(\mu_{i})$. The agent’s objective is to identify the arm with the highest $\mu_{i}$. Performance is measured using regret, calculated as $\sum_{t}(\mu_{a^{*}}-\mu_{\hat{a}_{t}})$. Unlike MDPs, MABs lack states.
In a Contextual MAB, each arm $a$ has a feature vector $x_{a}\in R^{d}$ (Sutton & Barto, 2018). At each step $t$, the agent observes a context state $s_{t}$. Selecting action $a_{i}$ results in a reward from a normal distribution with mean $\mu_{t}=\langle s_{t},a_{i}\rangle$ and standard deviation $\sigma$. Here, unlike in MDPs, the actions influence immediate rewards but not future states.
In-Context Learning. In-Context Learning describes the capability of a model to infer its task from the context it is given. For instance, the GPT-3 model (Brown et al., 2020) can be prompted in natural language to perform a variety of functions such as text classification, summarization, and translation, despite not being explicitly trained for these specific tasks. One of the possible prompts, which is used in our paper in some form, is a list of example pairs $(x_{i},y_{i})$ ending with a query $x_{q}$ for which the model is expected to generate a corresponding prediction $\hat{y}_{q}$.
Contrastive Learning. Contrastive learning focuses on creating representations where “similar” examples are close together in the feature space, while “dissimilar” ones are far apart (Weinberger & Saul, 2009; Schroff et al., 2015). This concept may be encapsulated in a triplet loss formula, where the similarity between an anchor and a positive example is maximized, and that between an anchor and a negative example is minimized, expressed as $L=sim(x,x^{+})-sim(x,x^{-})$. Oord et al. (2018) has developed a variant of contrastive loss called InfoNCE. This loss was lately widely adopted for representation learning (Jaiswal et al., 2020).
Appendix B Related Work
B.1 Transformers in Reinforcement Learning
According to the survey by Li et al. (2023), Transformers (Vaswani et al., 2017) are increasingly utilized in reinforcement learning (RL) for various tasks, including representation learning of individual observations and their histories, as well as model learning, as seen in Dreamer by Hafner et al. (2019). In our research, we focus on the application of Transformers in sequential decision-making and in developing generalist agents. Same as AD (Laskin et al., 2022), Headless-AD also utilizes transformers as the base model for our approach.
The incorporation of transformers as a sequence modeling tool in RL began with the Decision Transformer (DT) by (Hu et al., 2022), which trains autoregressively on offline datasets of state, action, and return-to-go tuples. Unlike conventional RL, which focuses on return maximization, DT generates appropriate actions during inference by conditioning on specified return-to-go values. The Trajectory Transformer by Janner et al. (2021) introduces an alternative approach using beam search to bias trajectory samples based on future cumulative rewards. Building on this, the Multi-Game Decision Transformer (MGDT) by Lee et al. (2022) improves upon DT with enhanced transfer learning capabilities for new games, eliminating the need for manual return specification. Similarly, the Switch Trajectory Transformer by (Lin et al., 2022) expands on the Trajectory Transformer to facilitate multi-task training. However, when transitioning to novel tasks, these approaches require fine-tuning the model.
B.2 Offline Meta-RL
Traditional RL agents, tailored to specific environments, struggle with novel tasks. In contrast, by training across a diverse array of tasks, Meta-RL equips agents with adaptable exploration strategies and an understanding of common environmental patterns (Duan et al., 2016; Wang et al., 2016; Rajeswaran et al., 2017; Zhang et al., 2018; Team et al., 2021; Nikulin et al., 2023). Offline Meta-RL, a subset of this approach, trains agents solely on pre-existing datasets, without direct environmental interaction. A critical challenge here is MDP ambiguity, where task-specific policies misinterpret data due to dataset biases. Dorfman et al. (2021) propose a data collection method to mitigate this issue, and suggest an approach that treats Meta-RL as a Bayesian RL problem for optimal exploration in new tasks. Li et al. (2020) introduce FOCAL, a framework separating task identification from control. However, their method relies on a strict mapping assumption that fails in certain scenarios, such as those with sparse rewards. Li et al. (2021) enhance FOCAL with an attention mechanism and improved metric learning, showing greater robustness in scenarios with sparse rewards and domain shifts. However, most Offline Meta-RL methods rely on explicit task modeling, which can introduce limiting biases. Alternatively, In-Context Learning implicitly infers tasks from environmental interactions, offering a potentially more flexible approach.
B.3 In-Context Learning in RL
In-Context Learning (ICL) is an ability of a pretrained model to adapt and perform a new task given a context with examples $\{x_{i},f(x_{i})\}^{n}$, where $f$ is a function that gives ground truth targets (Brown et al., 2020; Wang et al., 2023). Previous work on ICL (Mirchandani et al., 2023) shows that Large Language Models operate as General Pattern Machines and are able to complete, transform and improve token sequences, even when the sequences consist of randomly sampled tokens. This ability supports our design choice to randomly encode the actions of agents. The research on Transformer Circuits (Elhage et al., 2021) gives evidence of Transformers’ ability to copy tokens, either literally or on a more abstract level, explaining their ICL capabilities. This is particularly useful for our model, which explicitly incentivizes an abstract copying of the action embeddings.
Efforts to blend In-Context Learning (ICL) abilities of transformers with reinforcement learning (RL) are gaining traction, promising to create adaptable RL agents for real-world scenarios with varying conditions. Attempts to transfer ICL features to RL include Prompt-Based Decision Transformers (Xu et al., 2022), which leverage task-specific demonstration datasets as prompts, exhibiting strong few-shot learning without weight updates. Laskin et al. (2022) introduced Algorithm Distillation (AD) that trains a policy improvement operator using data from agent-environment interactions of learning RL algorithm, predicting the next action autoregressively. A key strategy here is across-episode training for capturing policy improvement. Lee et al. (2023) developed Decision Pretrained Transformer (DPT) using supervised training with unordered environmental interactions as context to predict optimal actions. Under certain conditions, DPT can be proved to implement posterior sampling, resulting in near-optimal exploration. However, DPT’s reliance on an optimal policy during training, not always available, is a limitation, similar to AD’s constraint on action space structure, restricting their use as fully generalist agents. Lin et al. (2023) provided theoretical analysis of AD and DPT, offering guarantees about learning from base algorithms. Zisman et al. (2023) propose a method for distilling the improvement operator from a demonstrator whose actions are initially noisy, with the noise level decreasing progressively throughout the data collection phase. This approach simplifies the data generation compared to AD as it does not require logging the training process of RL models and relaxes DPT’s limitation because the demonstrator may be suboptimal. Kirsch et al. (2023) demonstrated that data augmentation, through random projection of observations and actions, enhances task distribution and generalization on new domains, boosting ICL capabilities. Raparthy et al. (2023) study the effect of model size and dataset properties on the success of in-context learning. Other research, acknowledging Transformers’ limitations when it comes to long sequences, explores alternative models such as S4 (Gu et al., 2021). Lu et al. (2023) adapted S4 for RL, showing that it is capable of surpassing LSTM in performance and Transformers in runtime, indicating potential in RL ICL applications.
In contrast to previous works that tackle changing reward distributions or environment dynamics as changing goals, we expand the application of ICLRL to changing action spaces. Our use of random embeddings, motivated by the preparation of the model to unseen actions, may also improve ICL abilities by data augmentation, as suggested by Kirsch et al. (2023).
B.4 Discarding the Linear Layer
The Headless LLM approach, introduced by Godey et al. (2023), utilizes Contrastive Learning to eliminate the language head from the model architecture. Rather than generating a probability distribution over tokens, it directly predicts token embeddings. This strategy aims to enhance training and inference speeds by discarding a substantial linear layer. While in the RL context, where action spaces are typically small, this might not significantly impact runtime, it does offer a key advantage, as the model is no longer dependent on the number of actions and can therefore handle variable-length action spaces. Unlike Headless LLM, our approach does not rely on contrastive loss for learning action representations. Instead, we use fixed action embeddings, and the model is tasked to output the embedding of the next action. During inference, an action is chosen from a categorical distribution, where the logits are the similarities between the predicted embedding and available actions. Thus, the role of contrastive loss is to enhance the likelihood of selecting the correct action while diminishing the probabilities of other actions.
Wolpertinger by Dulac-Arnold et al. (2015) also employs a similar approach of removing the linear head to improve the train and inference speeds. They suggest associating each action with an embedding and performing the training in a continuous action space. A specific action index is chosen as a nearest neighbor of the predicted embedding, and its effect is treated as environment dynamics. Most importantly, this nearest neighbor selection process can be optimized using an approximate algorithm. This leads to logarithmic time complexity, which is particularly advantageous in environments with large action spaces. Thus, we believe that Headless-AD is also capable of performing well on large action spaces, benefiting from runtime gains of approximate nearest neighbor lookup. However, Wolpertinger’s usage of fixed action embeddings prevents introduction of new actions, highlighting the significance of Headless-AD’s usage of randomized action embeddings.
Appendix C Algorithm Distillation on Permuted Train Sets
Figure 10: AD Performance on Permuted Actions. This figure shows the models’ success rates on Darkroom environment, averaged over $5$ seeds. We evaluated AD’s ability to adapt to action sets with varied semantics by training it on a permuted dataset. Except for this, the training and testing are the same as in the vanilla setting and can be found in Section 4.3. Contrary to expectations, this tailored training did not enhance performance when compared to the AD model trained on standard datasets.
Appendix D MSE-Headless-AD In-Context Curves on Benroulli Bandit
Figure 11: MSE-Headless-AD In-Context Curves on Benroulli Bandit: Though Figure 8(b) shows that a variant of Headless-AD with MSE loss has a far from random regret, this graph shows that this result is due to the exploitation of several actions and the lack of exploration. We make this conclusion by observing that the shape of MSE-Headless-AD’s training curve is a straight line, showing the evidence that no learning is being present. All curves are averaged over $5$ seeds.
Appendix E Limitations and Future Work
Limitations. Headless-AD shares AD’s limitation of fixed sequence lengths, which may limit its effectiveness in environments with long episodes. A unique constraint of Headless-AD is its limit on the number of actions, dictated by the dimensionality of the action embeddings – only as many actions as there are dimensions can be orthogonally represented. Going over this limit causes embeddings to become linearly dependent, unintentionally distributing probability across multiple actions. While action spaces in RL environments typically do not become excessively large, and increasing the dimension of embeddings could mitigate this issue, it remains a point for consideration.
Future Work. In our study, we demonstrated the ability of our algorithm to generalize to new action spaces, as shown by its performance on elementary tasks. To extend and validate these findings, future research should focus on more complex environments. This will offer a deeper insight into the algorithm’s versatility and robustness in diverse and complex settings.
Additionally, to ensure the broader applicability and adaptability of our approach, it is essential to examine its compatibility and performance with various models beyond Algorithm Distillation (AD) (Laskin et al., 2022), such as the Decision Pretrained Transformer (DPT) (Lee et al., 2023). This exploration will provide a more comprehensive understanding of the algorithm’s strengths and limitations, potentially leading to further improvements and a wider scope of applications in different contexts.
Appendix F Sampling of Orthonormal Vectors
To sample the orthonormal vectors used as action embeddings, we use the torch.nn.init.orthogonal function from PyTorch (Paszke et al., 2019) that utilizes a specific algorithm from Saxe et al. (2013).
Appendix G Algorithms’ Training Times
Table 1: Algorithm Training Times. Here we list the training times (in hours) of Headless-AD and AD on each environment. All experiments were performed on A100 GPUs. Note that times here include both training and evaluation steps. Though Headless-AD requires more time for completion, we point out that it is evaluated on more environment sets compared to AD.
Headless-AD AD
Bernoulli Bandit $5.5$ $1.5$
Contextual Bandit $10$ $7$
Darkroom $27$ $19$
Appendix H Model Hyperparameters
In this section, we detail the hyperparameters for our models, each tuned for specific environments and design configurations. The tuning process utilized Bayesian optimization via the wandb sweep tool (Biewald, 2020). The optimization objective was chosen to maximize both the performance score achieved by the model and the efficiency in the number of actions utilized. The objective function was structured as follows:
[…]
where $n_{a}$ represents the total number of actions attempted by the model during evaluation, and $N$ signifies the amount of possible actions within an environment. Moreover, we took the return from the final episode in evaluation and normalized it to the $[0,1]$ range, with the lower bound 00 corresponding to the performance of a random agent, and the upper bound $1$ denoting the efficiency of our data generation algorithm. Darkroom environment already has the returns in the range $[0,1]$, so we did not perform normalization for this environment.
Table 2: Headless-AD’s Environment-Specific Hyperparameters: For certain instances, hyperparameters underwent optimization within the specified ranges in the Sweep Values column, utilizing the Bayesian search method facilitated by the wandb sweep tool (Biewald, 2020). This process was employed to identify the optimal set of hyperparameters for enhanced performance and fair comparisons.
Hyperparameter Bernoulli Bandit Darkroom Contextual Bandit Sweep Values
Number of Layers 4 4 8 [1,2,4,6,8]
Number of Heads 64 64 8 [2,4,8,16,32,64
Model Dim. 512 512 1024 [128, 256, 512, 1024, 2048]
Sequence Length 300 100 300
$\tau$ 1.34 7.8 3.0 [1e-5, 10]
Learning Rate 3.1e-04 1.7e-04 5.0e-05 [1e-5, 1e-2]
Weight Decay 4.4e-03 1.7e-02 5.0e-02 [1e-6, 5]
$\beta_{1}$ 0.68 0.72 0.5 [0, 1]
Attention Dropout Rate 0.22 0.7 0.3 [0, 1]
Dropout Rate 0.12 0.19 0.25 [0, 1]
In-Context Episodes 300 1000 300
Action Selection sample sample mode [sample, mode]
Table 3: AD’s Environment-Specific Hyperparameters: For certain instances, hyperparameters underwent optimization within the specified ranges in the Sweep Values column, utilizing the Bayesian search method facilitated by the wandb sweep tool (Biewald, 2020). This process was employed to identify the optimal set of hyperparameters for enhanced performance and fair comparisons.
Hyperparameter Bernoulli Bandit Darkroom Contextual Bandit Sweep Values
Number of Layers 4 2 8 [1,2,4,6,8]
Number of Heads 64 8 8 [2,4,8,16,32,64
Model Dim. 512 512 1024 [128, 256, 512, 1024, 2048]
Sequence Length 300 100 300
Label Smoothing 5.0e-03 5.0e-02 0.2 [0, 1]
Learning Rate 1.1e-05 4.0e-04 1.3e-05 [1e-5, 1e-2]
Weight Decay 1.5e-04 8.0e-06 0.25 [1e-6, 5]
$\beta_{1}$ 0.82 0.95 0.58 [0, 1]
Attention Dropout Rate 0.1 3.0e-02 0.28 [0, 1]
Dropout Rate 0.1 1.0e-02 0.45 [0, 1]
In-Context Episodes 300 1000 300
Action Selection sample mode mode [sample, mode]
Table 4: Headless-AD’s Environment-Specific Hyperparameters for Prompt Ablation: For certain instances, hyperparameters underwent optimization within the specified ranges in the Sweep Values column, utilizing the Bayesian search method facilitated by the wandb sweep tool (Biewald, 2020). This process was employed to identify the optimal set of hyperparameters for enhanced performance and fair comparisons.
Hyperparameter Bernoulli Bandit Darkroom Sweep Values
Number of Layers 4 4 [1,2,4,6,8]
Number of Heads 64 64 [2,4,8,16,32,64
Model Dim. 512 512 [128, 256, 512, 1024, 2048]
Sequence Length 300 100
Label Smoothing 0.92 8.5 [0, 1]
Learning Rate 5.0e-04 3.5e-04 [1e-5, 1e-2]
Weight Decay 2.6e-05 4.7e-03 [1e-6, 5]
$\beta_{1}$ 0.78 0.99 [0, 1]
Attention Dropout Rate 0.28 0.4 [0, 1]
Dropout Rate 0.16 0.16 [0, 1]
In-Context Episodes 300 1000
Action Selection sample sample [sample, mode]
Table 5: Headless-AD’s Environment-Specific Hyperparameters for Loss Ablation: For certain instances, hyperparameters underwent optimization within the specified ranges in the Sweep Values column, utilizing the Bayesian search method facilitated by the wandb sweep tool (Biewald, 2020). This process was employed to identify the optimal set of hyperparameters for enhanced performance and fair comparisons.
Hyperparameter Bernoulli Bandit Darkroom Sweep Values
Number of Layers 4 4 [1,2,4,6,8]
Number of Heads 64 64 [2,4,8,16,32,64
Model Dim. 512 512 [128, 256, 512, 1024, 2048]
Sequence Length 300 100
Learning Rate 5.0e-05 1.3e-05 [1e-5, 1e-2]
Weight Decay 0.17 2.5 [1e-6, 5]
$\beta_{1}$ 0.91 0.75 [0, 1]
Attention Dropout Rate 1.0e-03 8.0e-02 [0, 1]
Dropout Rate 0.14 0.75 [0, 1]
In-Context Episodes 300 1000
Action Selection mode mode
Appendix I Code Sample
from typing import Tuple
[…]
import torch.nn as nn
[…]
def __init__(
self,
action_embed_dim: int,
num_actions: int,
%**** icml2024.tex Line 950 ****device: str,
super().__init__()
self.action_embed_dim = action_embed_dim
self.num_actions = num_actions
self.device = device
[…]
self.action_map.requires_grad = False
[…]
def regenerate(self) -> None:
torch.nn.init.orthogonal_(self.action_map, gain=1)
[…]
def __call__(self, actions: torch.Tensor) -> torch.Tensor:
embeds = self.action_map[actions]
[…]
def get_action(
self,
embeds: torch.Tensor,
) -> Tuple[torch.Tensor, torch.Tensor]:
%**** icml2024.tex Line 975 ****sims = embeds @ self.action_map.T
[…]
actions_sample, actions_argmax = dist.sample(), dist.probs.argmax(-1)
[…]
def _get_action_map_as_context(self, batch_size: int):
actions = torch.tile(self.action_map.unsqueeze(0), (batch_size, 1, 1))
[…]
act_mapper.regenerate()
action_embeds = act_mapper(actions)
[…]
states=states,
act_emb=action_embeds,
%**** icml2024.tex Line 1000 ****rewards=rewards,
actions_list=actions_list,
[…]
Figure 12: Code that demonstrates the Headless-AD training procedure. Note that this snippet is intended for illustration purposes only. The complete code can be found in Headless-AD’s repository, which will be open-sourced after the review.